Calculate the pH when 0.020 mole of HBr is added to 200.0 mL of a solution containing 0.60 M NH3 and 0.60 M NH4Cl, assuming no volume change. (Kb of NH3 =1.8x10-5). Report your answer to three significant figures.
1
Expert's answer
2020-03-25T12:47:53-0400

"pOH = pKb + log(\frac{salt}{base})."

"pH + pOH = 14."

Kb= 1.8*10-5

pOH=(1.8*10-5)+log(0.6/0.1)= 6.00

Therefore; pH=14-6 =8